OpenStudy (anonymous):

Solve the exponential equation for x. 18 + 5x = 43

OpenStudy (anonymous):

not really exponential is it?

OpenStudy (anonymous):

\[18+5^x=43\]

OpenStudy (anonymous):

\[5^x=43-18=25\] si \[x=2\] since \[5^2=25\]

OpenStudy (saifoo.khan):

25